OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 16 less than or greater to 11 and x + 16 greater than or equal to 3 11 greater than or equal to x + 16 greater than or equal to 3 -5 greater than or equal to x greater than or equal to -13 A number line with a closed circle on -13, a closed circle on -5, and shading in between. A number line with a closed circle on -13, shading to the left, and a closed circle on -5, shading to the right.

OpenStudy (anonymous):

http://www.chegg.com/homework-help/definitions/compound-inequalities-27 check it out and let me know if you have any questions

OpenStudy (anonymous):

I don't get it..

OpenStudy (anonymous):

create a graph on the number line for each statement and you see that one of these isn't a compound inequality.

OpenStudy (anonymous):

there is no such thing as "less than or greater to"

OpenStudy (anonymous):

I cant open, the attachment

OpenStudy (anonymous):

@satellite73 those were the answers I got, if that's the question you're asking.

OpenStudy (texaschic101):

x + 6 <= 11 x <= 11 - 6 x <= 5 x + 16 >= 3 x >= 3 - 16 x > = - 13 The last one is incorrect

OpenStudy (texaschic101):

any questions ?

OpenStudy (texaschic101):

your last answer choice is saying a number line with a closed circle on -13(this part is true), however, it also says shading to the left.(wrong) less then means shading to the left, and greater then means shading to the right. The answer is x >= -13, that says greater, meaning shading to the right, not the left.

OpenStudy (anonymous):

that helped a lot, thank you @texaschic101

OpenStudy (texaschic101):

your quite welcome :)